OpenStudy (anonymous):

1/27 to the power of negative two-thirds?

OpenStudy (anonymous):

1/27 to the power of negative two-thirds = 27 to the power of positive two-thirds or 27^2/3 which equals (27^2) / 3 which equals 243.

OpenStudy (kira_yamato):

27^2/3 = 3^2 = 9